OpenStudy (anonymous):

find the slope of the line containing points with coordinates (-3,7) and (3,2) 0 - 5/6 undefined 5/6

OpenStudy (katrinakaif):

Didn't we do similar problems to these yesterday? Your problem has given points: (-3,7) : (x1,y1) (3,2) : (x2,y2) The formula for finding slope is: (y2-y1) ------- (x2-x1) Simply substitute: (2 - 7) ----- (3 - (-3) and solve

OpenStudy (anonymous):

slope m = ( 7 -2) / ( -3 -3) = -5/6 Vince, Why don't you start practicing on slope topic? I don't mind solving these, but it'd be more benefit for you to do it yourself!